OpenStudy (ccastudentgrade10):

fan & medal :) solve by completing the square? -x^2 + 6x + 10 = 0 it's bonus for math my teacher said even though we havent learned it if you do it you get two points added to the exam and i could really use help! ps you have to show your work

OpenStudy (ccastudentgrade10):

@aloud ?

OpenStudy (dustinrathke24):

-(x+3)(x+3)=10 -(x+3)^2=10 -x+3=+- square root of 10=3.162 -x=3.162-3 or -x=-3.162-3 x=-.162 or x=6.162

OpenStudy (ccastudentgrade10):

thank you!!!!!